# Machine Learning Model Integration with Open World Temporal Logic for Process Automation

### v1

@@ 4.2.1 Logical Encoding in PyReason
We encode game state using five predicates and three rules:
@@
  • hand_as_point_vals(hand): This rule updates the cumulative point value of the hand by appending each drawn card’s value. It uses a non-grounded rule with an annotation function (append_hand_annotation_fn) to incrementally build the hand score. The threshold $[0.3,1]$ ensures triggering for any card.
@@
Figure 5 depicts the full simulation loop: one thread monitors termination conditions, while another periodically classifies drawn cards and updates the logical state. This concurrent design shows how PyReason’s fixed-point inference can be driven byand synchronized with–real (wall-clock) time. The classification and inference module periodically polls the image classifier to identify each drawn card and then encodes those results as PyReason facts, enabling logical reasoning about the evolving game in the background.
@@ 4.2.2 Interpreting the Inference Trace
    Asserts each drawn card via the card_drawn_obj fact at certainty $[1,1]$.
@@
    Fires odds_of_losing to compute the interval $[p,1]$, where $p$ is the fraction of remaining cards that would push the total above 42. When $[p,1]=[1,1]$, no safe draws remain and the game ends.
@@ 5 Conclusions
The framework’s utility was demonstrated in two case studies: a welding scenario for real-time process control and a card game for managing uncertainty in decision-making. These examples highlighted the system’s ability to handle temporal dynamics and produce fully explainable inference traces, which are crucial for debugging and policy refinement.
Future work will focus on extending this integration in several key directions. In particular, we look to integrate “metacognitive” learned about the failure of the model to improve results on down-stream tasks (e.g., see [9]). In particular, early results have shown the combination of such rules with abductive inference in PyReason can lead to improved ensembling of object detection ML models [10]. Based on those results, we look to examine further practical integrations for ML-logic programming integration with such error correciton capabilities. Additionally, we plan to apply the framework to more complex, large-scale industrial scenarios and explore its use with a broader range of ML models, such as Large Language Models for business process automation, to further validate its robustness and versatility.

### v2

@@ 4.2.1 Logical Encoding in PyReason
We encode the game state using five predicates and three rules:
@@
  • hand_as_point_vals(hand): This rule updates the cumulative point value of the hand by appending each drawn card’s value. It uses a non-grounded rule with an annotation function (append_hand_annotation_fn) to incrementally build the hand score. The threshold [0.3,1][0.3,1][ 0.3 , 1 ] ensures triggering for any card.
@@
Figure 5 depicts the full simulation loop: one thread monitors termination conditions, while another periodically classifies drawn cards and updates the logical state. This concurrent design shows how PyReason’s fixed-point inference can be driven by — and synchronized with — real (wall-clock) time. The classification and inference module periodically polls the image classifier to identify each drawn card and then encodes those results as PyReason facts, allowing for logical reasoning about the evolving game in the background.
@@ 4.2.2 Interpreting the Inference Trace
    Asserts each drawn card via the card_drawn_obj fact at certainty [1,1][1,1][ 1 , 1 ].
@@
[…]
@@ 5 Conclusions
The utility of the framework was demonstrated in two case studies: a welding scenario for real-time process control and a card game for managing uncertainty in decision making. These examples highlighted the system’s ability to handle temporal dynamics and produce fully explainable inference traces, which are crucial for debugging and policy refinement.
Beyond the integration with machine learning models for perception, the PyReason framework is architecturally designed to support the integration of external, specialized reasoning engines, a concept analogous to the pluggable constraint solvers found in Constraint Logic Programming (CLP). This modularity allows the logic program to function as a high-level orchestrator that defines a problem space and coordinates with specialized computational modules. While the core of this paper focuses on integrating machine learning models for perception, the framework’s underlying design is more general, enabling interfaces with various external solvers. A clear demonstration of this capability is presented in the work [12], which describes using PyReason to manage medical triage optimization. In their approach, the logic program dynamically formulates optimization problems by reasoning over situational facts to select the appropriate objectives and constraints. These problem instances are then passed to a dedicated external integer programming solver for resolution. Crucially, the solver’s output is translated back into logical facts and reintegrated into the PyReason program, which then performs further symbolic reasoning on these results to guide decision-making. This workflow exemplifies a general pattern where PyReason orchestrates a process by leveraging an external, “pluggable” engine for a specialized reasoning task, while maintaining a single, explainable logical framework.
Future work will focus on extending this integration in several key directions. In particular, we look to integrate “metacognitive” learned about the failure of the model to improve results on down-stream tasks (e.g., see [9]). In particular, early results have shown that the combination of such rules with abductive inference in PyReason can lead to improved ensembling of object detection ML models [10]. Based on those results, we look to examine further practical integrations for ML-logic programming integration with such error correction capabilities. Additionally, we plan to apply the framework to more complex, large-scale industrial scenarios and explore its use with a broader range of ML models, such as Large Language Models for business process automation, to further validate its robustness and versatility.